OpenStudy (anonymous):

Write an equation in slope intercept form then in standard form. m=3 and the y-intercept is (0,2)

OpenStudy (kohai):

What is point slope form?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (kohai):

Right. So your points are going to go in x1 and y1 and your slope is going to be m. What will that look like?

OpenStudy (anonymous):

y-2=3(x-0) that comes down to y=3x+2 right?

OpenStudy (jdoe0001):

yeap

OpenStudy (kohai):

Yes

OpenStudy (anonymous):

so how do i put that in standard form now?

OpenStudy (anonymous):

standard form is Ax+by=c

OpenStudy (anonymous):

o i got it! is it -2=3x-y?